two planes leave the same point at the same time. One plane travels at 252mi/h. The other plane travels at 266mi/h. How long will it take befor they are 1750 miles apart?
Answer by checkley77(12844)

YOU NEGLECTED TO MENTION WHICH DIRECTION EACH PLANE IS TRAVELING SO I'LL CALCULATE THE SAME DIRECTION & THEN OPPOSITE DIRECTIONS.
RT=D
266H-252H=1,750
14H=1,750
H=1,750/14
H=125 HOURS.
266H+252H=1,750
518H=1,750
H=1,750/518
H=3.38 HOURS.